# HGAN-SDEs: Learning Neural Stochastic Differential Equations with Hermite-Guided Adversarial Training

### v1

@@ 1 Introduction
Stochastic Differential Equations (SDEs) [ito1944, stratonovich1964new, ocone1988stochastic] are widely used to model dynamic systems under random fluctuations. However, they rely on predefined drift and diffusion functions, which are rarely observable and cannot be specified in advance. This mismatch limits their applicability in real-world modeling. Neural stochastic differential equations (Neural SDEs) [tzen2019neural] address this issue by integrating neural networks with SDEs, enabling drift and diffusion dynamics to be learned from data.
Training Neural SDEs is notoriously difficult because exact likelihoods are intractable. GANs [kidger2021neural] provide a likelihood-free alternative by matching trajectory distributions through a parametric discriminator. Prior work has employed CDE-based discriminators, but such approaches are often unstable and computationally expensive, limiting scalability. We therefore seek a discriminator design that remains likelihood-free while incorporating tractable approximations of SDE dynamics. To this end, we propose a novel GAN framework based on neural Hermite functions. Hermite functions approximate SDE transition probabilities, enabling the discriminator to effectively assess whether a generated path aligns with the underlying DGP. In addition, their statistical properties enhance optimization stability and computational efficiency.
@@ 2 Related Work
Neural stochastic differential equations (Neural SDEs) extend neural ordinary differential equations (Neural ODEs) [chen2018neural] by adding stochastic diffusion terms, which improve long-horizon stability and enable modeling of random perturbations [tzen2019neural, Liu2019neural]. They have since been applied in domains such as finance [cuchiero2020generative], scientific machine learning [rackauckas2020universal], manifold learning [zeng2024latent], and generative modeling [song2020score]. Training, however, remains challenging: stochasticity in both the dynamics and gradient estimates hinders convergence and increases computational cost [li2020scalable]. Existing approaches address these challenges along two main directions: non-adversarial (variational or kernel-based) and adversarial (GAN-based). Non-adversarial methods target either endpoint distributions or entire trajectories, employing reparameterized variational objectives [li2020scalable], signature kernels [issa2024non], or low-variance gradient estimators [course2024amortized, xu2022infinitely]. Adversarial methods instead match real and generated paths via discriminators, and have shown promise in capturing complex or heavy-tailed dynamics, particularly in financial applications [cuchiero2020generative, Gierjatowicz2020robust]. Nevertheless, non-adversarial methods often rely on restrictive approximations, while adversarial ones suffer from instability and high computational cost—highlighting the need for more efficient training frameworks.
@@ 3.2 Neural Hermite Functions as GAN Discriminators
To approximate this stationary density $p(x)$, we propose using Hermite functions as a basis expansion: $p(x)\approx\sum_{n=0}^{N}c_{n}\psi_{n}(x),$ where the $\psi_{n}(x)$ are the Hermite functions defined by: $\psi_{n}(x)=\frac{1}{\sqrt{2^{n}n!\sqrt{\pi}}}e^{x^{2}/2}(-1)^{n}\frac{d^{n}}{dx^{n}}e^{-x^{2}}.$ Hermite functions enjoy a critical property: they form a complete orthonormal basis in the Hilbert space $L^{2}(\mathbb{R},dx)$. This property guarantees that any square-integrable function—including valid probability densities—can be approximated arbitrarily well by a linear combination of these functions. We formalize this in the following theorem:
Theorem 1 (Orthonormality of Hermite functions [abramowitz1972handbook, eijndhoven1990new]).
@@
In this work, we approximate the stationary transition density of an SDE as a truncated Hermite expansion up to order $N$, formulated as:
@@
Following [eijndhoven1990new], integrating both sides of (1) leads to an identity involving a constant $C$. A further integration over the real line shows that the integrals remain finite, which forces $C=0$. Hence we obtain the reduced form
@@
3.3 Expressivity of Hermite-Based Discriminator
Theorem 3 (Discriminative Power of Hermite Span).
If Assumption 1 holds and the true transition density $p(x(t)\mid x(0)=x_{0},t)$ is square-integrable, then
$d_{\mathcal{P}}(\mu,\nu)=0\Longleftrightarrow\mu=\nu.$
Proof.
According to [stekloff1916theoreme], Hermite polynomials $\{H_{n}(x)\}$ are dense in $L^{2}(\mathbb{R},e^{-x^{2}}dx)$, implying that $\{\psi_{n}(x)\}$ are dense in $L^{2}(\mathbb{R},dx)$. By Theorem 2.2 of [DBLP:conf/iclr/Zhang0ZX018], the induced metric $d_{\mathcal{P}}(\mu,\nu)$ is therefore discriminative. ∎
@@ 3.3 Expressivity of Hermite-Based Discriminator
Linear Model DLinear [zeng2022dlinear] 6.00 3.20 0.09 0.11 61.61 9.09 0.52 0.31 196.79 7.01 1.34 0.44 5.18 3.86 0.08 0.13
Convolutional Models MICN [wang2023micn] 7.05 4.74 0.12 0.16 159.01 9.55 1.97 0.41 33.74 6.87 0.59 0.23 7.65 4.48 0.11 0.12
TimesNet [wu2023timesnet] 1.50 0.21 0.03 0.02 13.56 8.10 0.14 0.15 2.42 2.70 0.05 0.07 1.86 1.74 0.04 0.04
Attention-based Models Transformer [vaswani2017attention] 0.74 0.55 0.02 0.03 1.22 0.70 0.03 0.02 1.44 0.93 0.03 0.04 9.47 10.00 0.23 0.40
Informer [zhou2021informer] 3.53 3.49 0.05 0.10 8.58 3.20 0.10 0.09 1.57 2.46 0.04 0.07 67.67 7.13 0.99 0.41
Autoformer [wu2021autoformer] 1.71 1.53 0.03 0.06 10.87 9.22 0.13 0.15 2.39 4.41 0.05 0.08 145.31 6.33 1.05 0.33
RNN Model SegRNN [wang2023segrnn] 1.09 0.31 0.02 0.02 13.68 7.24 0.16 0.14 2.12 2.11 0.04 0.06 4.98 7.55 0.09 0.28
State Space Model Mamba [gu2023mamba] 1.04 0.10 0.02 0.01 13.72 7.58 0.17 0.14 2.39 2.55 0.05 0.07 1.35 2.21 0.03 0.06
SDE-based Models Latent-SDEs [rubanova2019latent] 0.44 10.00 0.01 0.42 101.40 10.00 2.51 0.93 39.78 10.00 0.98 0.86 3.73 6.81 0.06 0.20
GAN-SDEs [kidger2021neural] 0.38 9.27 0.01 0.32 214.54 10.00 4.11 0.88 209.23 10.00 3.91 0.87 0.11 0.27 0.02 0.02
@@
4 Experiments
4.1 Experiment Settings
Benchmarks. Following Yacine’s work [ait1999transition], we construct four simulation datasets to evaluate our framework: geometric Brownian motion (GBM), Ornstein-Uhlenbeck (OU), Cox-Ingersoll-Ross (CIR), and Polynomial Drift. Their parameter configurations are summarized in Table 2. Each dataset consists of 20,000 training and 6,000 test samples, formatted for sequence prediction: the first 100 points are given as inputs and the subsequent 50 as prediction targets. For real-world validation, we use three datasets. Stock-AAL and Stock-ADBE each contain 43,146 five-minute stock prices (2017–2018) with several hundred missing values, reflecting the stochastic volatility of financial markets. The Traffic dataset comprises 21,600 five-minute records of Beijing traffic flows in 2022, with no missing values and strong periodicity.
@@ 5 Conclusion
Acknowledge
Guang Zhang acknowledges the support from the Guangzhou-HKUST(GZ) Joint Funding Program (No.2024A03J0630).

### v2

@@ 1 Introduction
Random fluctuations are ubiquitous in complex dynamical systems across fields such as finance [27, 28], physics [16], and biology [6]. To capture these dynamics, Stochastic Differential Equations (SDEs) [10, 18] serve as a powerful tool for rigorous modeling and analysis. However, they rely on predefined drift and diffusion functions, which are rarely observable and cannot be specified in advance. This mismatch limits their applicability in real-world modeling. Neural stochastic differential equations (Neural SDEs) [19] address this issue by integrating neural networks with SDEs, enabling drift and diffusion dynamics to be learned from data.
Training Neural SDEs is notoriously difficult because exact likelihoods are intractable. GANs [11] provide a likelihood-free alternative by matching trajectory distributions through a parametric discriminator. Prior work has employed CDE-based discriminators, but such approaches are often unstable and computationally expensive, limiting scalability. We therefore seek a discriminator design that remains likelihood-free while incorporating tractable approximations of SDE dynamics. To this end, we propose a novel GAN framework based on neural Hermite functions. Hermite functions approximate SDE transition probabilities, enabling the discriminator to effectively assess whether a generated path aligns with the underlying DGP. In addition, their statistical properties enhance optimization stability and computational efficiency.
@@ 2 Related Work
Neural stochastic differential equations (Neural SDEs) extend neural ordinary differential equations (Neural ODEs) [3] by adding stochastic diffusion terms, which improve long-horizon stability and enable modeling of random perturbations [19, 13]. They have since been applied in domains such as finance [5], scientific machine learning [14], manifold learning [30], and generative modeling [17]. Training, however, remains challenging: stochasticity in both the dynamics and gradient estimates hinders convergence and increases computational cost [12]. Existing approaches address these challenges along two main directions: non-adversarial (variational or kernel-based) and adversarial (GAN-based). Non-adversarial methods target either endpoint distributions or entire trajectories, employing reparameterized variational objectives [12], signature kernels [9], or low-variance gradient estimators [4, 26]. Adversarial methods instead match real and generated paths via discriminators, and have shown promise in capturing complex or heavy-tailed dynamics, particularly in financial applications [5, 7]. Nevertheless, non-adversarial methods often rely on restrictive approximations, while adversarial ones suffer from instability and high computational cost—highlighting the need for more efficient training frameworks.
@@ 3.2 Neural Hermite Functions as GAN Discriminators
To approximate this stationary density $p(x)$, we propose using Hermite functions as a basis expansion: $p(x)\approx\sum_{n=0}^{N}c_{n}\psi_{n}(x),$ where the $\psi_{n}(x)$ are the Hermite functions defined by:
$\psi_{n}(x)=\frac{1}{\sqrt{2^{n}n!\sqrt{\pi}}}e^{x^{2}/2}(-1)^{n}\frac{d^{n}}{dx^{n}}e^{-x^{2}}.$
Hermite functions enjoy a critical property: they form a complete orthonormal basis in the Hilbert space $L^{2}(\mathbb{R},dx)$. This property guarantees that any square-integrable function—including valid probability densities—can be approximated arbitrarily well by a linear combination of these functions. We formalize this in the following theorem:
Theorem 1 (Orthonormality of Hermite functions [1, 20]).
@@
In this work, we approximate the stationary transition density of an SDE as a truncated Hermite expansion of order $N$, formulated as:
@@
Following [20], integrating both sides of (1) leads to an identity involving a constant $C$. A further integration over the real line shows that the integrals remain finite, which forces $C=0$. Hence we obtain the reduced form
@@
Linear Model DLinear [29] 6.00 3.20 0.09 0.11 61.61 9.09 0.52 0.31 196.79 7.01 1.34 0.44 5.18 3.86 0.08 0.13
Convolutional Models MICN [22] 7.05 4.74 0.12 0.16 159.01 9.55 1.97 0.41 33.74 6.87 0.59 0.23 7.65 4.48 0.11 0.12
TimesNet [24] 1.50 0.21 0.03 0.02 13.56 8.10 0.14 0.15 2.42 2.70 0.05 0.07 1.86 1.74 0.04 0.04
Attention-based Models Transformer [21] 0.74 0.55 0.02 0.03 1.22 0.70 0.03 0.02 1.44 0.93 0.03 0.04 9.47 10.00 0.23 0.40
Informer [32] 3.53 3.49 0.05 0.10 8.58 3.20 0.10 0.09 1.57 2.46 0.04 0.07 67.67 7.13 0.99 0.41
Autoformer [25] 1.71 1.53 0.03 0.06 10.87 9.22 0.13 0.15 2.39 4.41 0.05 0.08 145.31 6.33 1.05 0.33
RNN Model SegRNN [23] 1.09 0.31 0.02 0.02 13.68 7.24 0.16 0.14 2.12 2.11 0.04 0.06 4.98 7.55 0.09 0.28
State Space Model Mamba [8] 1.04 0.10 0.02 0.01 13.72 7.58 0.17 0.14 2.39 2.55 0.05 0.07 1.35 2.21 0.03 0.06
SDE-based Models Latent-SDEs [15] 0.44 10.00 0.01 0.42 101.40 10.00 2.51 0.93 39.78 10.00 0.98 0.86 3.73 6.81 0.06 0.20
GAN-SDEs [11] 0.38 9.27 0.01 0.32 214.54 10.00 4.11 0.88 209.23 10.00 3.91 0.87 0.11 0.27 0.02 0.02
@@
3.3 Expressivity of Hermite-Based Discriminator
Theorem 3 (Discriminative Power of Hermite Span).
If Assumption 1 holds and the true transition density $p(x(t)\mid x(0)=x_{0},t)$ is square-integrable, then $d_{\mathcal{P}}(\mu,\nu)=0\Longleftrightarrow\mu=\nu.$
Proof.
According to [stekloff1916théorème], Hermite polynomials $\{H_{n}(x)\}$ are dense in $L^{2}(\mathbb{R},e^{-x^{2}}dx)$, implying that $\{\psi_{n}(x)\}$ are dense in $L^{2}(\mathbb{R},dx)$. By Theorem 2.2 of [31], the induced metric $d_{\mathcal{P}}(\mu,\nu)$ is therefore discriminative. ∎
4 Experiments
4.1 Experiment Settings
Benchmarks. Following Yacine’s work [2], we construct four simulation datasets to evaluate our framework: geometric Brownian motion (GBM), Ornstein-Uhlenbeck (OU), Cox-Ingersoll-Ross (CIR), and Polynomial Drift. Their parameter configurations are summarized in Table 2. Each dataset consists of 20,000 training and 6,000 test samples, formatted for sequence prediction: the first 100 points are given as inputs and the subsequent 50 as prediction targets. For real-world validation, we use three datasets. Stock-AAL and Stock-ADBE each contain 43,146 five-minute stock prices (2017–2018) with several hundred missing values, reflecting the stochastic volatility of financial markets. The Traffic dataset comprises 21,600 five-minute records of Beijing traffic flows in 2022, with no missing values and strong periodicity.